OpenStudy (anonymous):

find the x intercept of y =4/5(x-15)

OpenStudy (anonymous):

\[\frac{4}{5}(x-15) = 0\]\[(x-15) = 0\]\[x = 15\] It intercepts the x-axis, so the y is evidently 0: Therefore the answer is (15, 0)